A magnet is suspended so that it may oscillate in the horizontal plane .It performs oscillations per minute at a placewhere  the angle of dip is 300  and 15 oscillations per minute, where the angle of dip is 600. The ratio of earth’s magnetic field at these two places is
1) 16/9              2)4/3             3)8/3        4)16/9

Grade:12th pass

1 Answers

ROSHAN MUJEEB
3 years ago
Given v1=2060=13sec−1 and v2=1560=14sec−1
Now v=12πMBHI−−−−−−√=12πMBcosφI−−−−−−−−√(∵BH=Bcosφ)
∴v1v2=B1cosφ1B2cosφ2−−−−−−−−√⇒B1B2=(v1v2)2(cosφ2cosφ1)2
B1B2=(1/31/4)2cos60∘cos30∘=169×1/23–√/2=1693–√